# BECoTTA: Input-dependent Online Blending of Experts for Continual Test-time Adaptation

### v2

@@ 3.3 Continual Test-time Adaptation Process
Model Initialization. We initialize our BECoTTA and baselines into two different settings: random initialization and post-pre-training with a Weather-Augmented Dataset (WAD). Following recent trends in CTTA (Song et al., 2023; Niu et al., 2022; Lim et al., 2023; Choi et al., 2022a; Liu et al., 2021; Adachi et al., 2022; Jung et al., 2023) and for a fair comparison, we perform a short pre-training for trainable parameters in models before deploying them to the CTTA problems. (Please refer to Tab. 1). For the pre-training initialization, we first build the WAD with $D$ domains. Next, we introduce a Domain Discriminator (DD) as the auxiliary head. It consists of lightweight CNN layers and is trained on WAD to classify the pre-defined $D$ domains. This helps the model distinguish between different domains and classify each test-time image input accordingly during test-time adaptation on a sequence of unseen domains. Then, we update MoDE layers and DD for only a small number of epochs. The total initialization loss $L_{init}$ is formulated below. Except for original cross-entropy loss $L_{seg}$ for semantic segmentation, we include cross-entropy loss $L_{disc}$ for DD, and domain-expert synergy loss $\Theta(D;A)$:
@@
Source-free TTA with MoDE. Building upon the initialized MoDE, we deploy our BECoTTA to real-world continual scenarios. Note that we do not access any source dataset after the deployment, maintaining the source-free manner in the test time as other prior works. In the source-free TTA, we only activated MoDE layers to transfer the target domain knowledge efficiently. Utilizing the frozen DD trained at initialization, we get the pseudo domain label $d$ per each domain-agnostic target image $x_{t}^{c}$. Afterward, according to $d$, we initially assign the domain-wise router and proceed with the aggregation of domain-adaptive experts. This approach ensures that our BECoTTA maintains input dependency, even within a dynamic continual scenario.
Following the preliminary works (Wang et al., 2020; Song et al., 2023; Niu et al., 2022), we adopt the entropy minimization using $H(\hat{y_{c}})=-\sum p(\hat{y_{c}})\cdot\log p(\hat{y_{c}})$. To prevent forgetting and error accumulation, we conduct entropy filtering based on the confidence of the pseudo labels. Therefore, the entropy-based loss $L_{tta}$ is as follows:
@@ 4 Experiments
4.1 Data Setup
Round 1 10 $\Delta$ Parameter
w/o WAD SAR (Niu et al., 2023) - 41.0 64.5 33.4 54.5 46.6 48.0 41.3 64.3 31.6 54.2 46.6 47.6 -0.4 0.02M
BECoTTA (Ours)-S MoDE 42.9 68.5 35.0 57.2 47.8 50.5 43.0 68.5 35.1 57.3 48.8 50.7 +0.1 0.09M
BECoTTA (Ours)-M MoDE 43.8 68.8 34.9 57.9 49.2 50.9 43.7 68.8 34.5 57.9 49.2 50.9 +0.0 0.63M
BECoTTA (Ours)-L MoDE 43.9 69.1 35.0 58.3 50.2 51.3 44.0 69.1 35.1 58.3 50.2 51.3 +0.0 3.16M
w/ WAD SAR (Niu et al., 2023) Full 43.6 68.6 44.5 59.1 48.7 52.9 43.4 67.4 42.2 58.1 47.6 51.9 -1.0 0.02M
BECoTTA (Ours)-S MoDE 44.1 69.5 40.1 56.8 49.1 51.9 44.0 69.4 40.1 56.9 49.1 51.9 +0.0 0.12M
BECoTTA (Ours)-M MoDE 45.6 70.8 42.6 59.6 50.8 53.9 45.6 70.7 42.5 59.5 50.8 53.9 +0.0 0.77M
BECoTTA (Ours)-L MoDE 45.7 71.4 43.7 59.6 50.5 54.2 45.7 71.3 43.7 59.6 50.6 54.2 +0.0 3.32M
@@ 4.1 Data Setup
Continual Disjoint Shifts (CDS) benchmark. To reflect various domain shifts, we adopt balanced weather shifts (CDS-Easy) and imbalanced weather & area shifts (CDS-Hard) scenarios. For the CDS-Easy, we utilize the Cityscapes-ACDC setting used in previous work (Wang et al., 2022a): Cityscapes (Cordts et al., 2016) is set to the source domain, and ACDC (Sakaridis et al., 2021) is set to the target domain, consisting of four different weather types (fog, night, rain, snow) with a same 400 number of instances. For the CDS-Hard, we propose a new imbalanced scenario both considering weather and geographical domain shifts. We additionally add clear and overcast weather from BDD-100k (Yu et al., 2020) to the existing target domain to mimic the real-world variety.
@@
Continual Gradual Shifts (CGS) benchmark. To construct gradually changing weather scenarios with blurry boundaries, we first conduct sampling from a Gaussian distribution with CDS-Easy target domains. Given a total of 1600 (400x4) timesteps in one round (including four tasks) at CDS-Easy, we define sampling distributions $\bm{\theta_{i}}\sim\mathcal{N}(400i,200)$ for each domain $i$, and perform uniform sampling to represent gradual changes of weathers. In the end, we construct four tasks containing blurry boundaries of weather as illustrated in Fig. 3.
@@
Domain Generalization (DG) benchmark. To demonstrate the versatility of our method, we conduct additional zero-shot experiments using the DG benchmark (Choi et al., 2021). This benchmark includes two large-scale real-world data sets (BDD-100k (Yu et al., 2020), Mapillary (Neuhold et al., 2017)) and two simulated data sets (GTAV (Richter et al., 2016), Synthia (Ros et al., 2016)).
4.2 Experimental Setting
Baselines. We compare our model with strong continual test-time adaptation methods including TENT (Wang et al., 2020), CoTTA (Wang et al., 2022a), SAR (Niu et al., 2023), EcoTTA (Song et al., 2023), VDP (Gan et al., 2023), DePT (Gao et al., 2022b), TTN (Lim et al., 2023). More details are found in Appendix (Sec. A).
Evaluation metric. All of the semantic segmentation results are reported mIoU in %. For the overall scenarios, we repeat each task in 10 rounds (a few rounds are reported for visibility). Please refer to the Appendix (Sec. C) for detailed results.
Implementation details of BECoTTA. Our BECoTTA has a flexible design, so it provides multiple variants according to the selection of rank and the number of experts. As such, we categorize the results into three groups: S, M, and L. We adopt four experts for S, and only injected MoDE into the last block of the encoder. Both M and L utilize six experts, with the only difference being the rank. For the CDS-Easy scenario, we leverage the pre-trained Segformer-B5 as our source model, aligning with CoTTA (Wang et al., 2022a). For other scenarios, we opt for Segformer-B2. Note that there is a difference between the two setups due to the size of Segformer, but we unify the Ours-S, M, and L architecture settings for all scenarios. To implement the initialization process, we warm up our architecture for 10 epochs like previous works (Song et al., 2023; Lim et al., 2023). Other concrete implementation details are found in Appendix (Sec. B).
Fairness with other baselines. For a fair comparison, we report both w/o WAD and w/WAD results for all experiments. w/o WAD means WAD is not used for initialization and it is directly compared with other baselines. We are fully aware that other baselines (Wang et al., 2022a, 2020; Niu et al., 2023) do not need to initialize, however, we conduct the same slight initialization while activating full parameters of source model to measure the effectiveness of WAD on them. Nevertheless, if only a part of the model activates, such as EcoTTA and ours, the rest of the model is frozen, and only that specific part is trained with WAD. In conclusion, the parameter difference between w/WAD and w/o WAD settings arises from the presence of domain-wise routers.
@@ 4.2 Experimental Setting
BN Stats Adapt (Nado et al., 2020) - 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 52.0 0.09M
@@
BECoTTA (Ours)-S 71.3 41.1 62.4 59.8 71.3 41.1 62.4 59.8 71.4 41.1 62.4 59.8 58.6 0.09M
                                   + WAD 72.0 45.4 63.7 60.0 71.7 45.2 63.6 60.1 71.7 45.4 63.6 60.1 60.2 0.12M
BECoTTA (Ours)-M 72.3 42.0 63.5 60.1 72.4 41.9 63.5 60.2 72.3 41.9 63.6 60.2 59.5 2.15M
                                   + WAD 71.8 48.0 66.3 62.0 71.7 47.7 66.3 61.7 71.8 47.7 66.3 61.9 61.9 2.70M
BECoTTA (Ours)-L 71.5 42.6 63.2 59.1 71.5 42.6 63.2 59.1 71.5 42.5 63.2 59.1 59.1 11.31M
                                   + WAD 72.7 49.5 66.3 63.1 72.6 49.4 66.3 62.8 72.5 49.7 66.2 63.1 63.0 11.86M
4.3 Main Results
CDS-Hard (imbalanced weather & area shifts). As shown in Tab. 1 and Fig. 1, all of our BECoTTA-S/M/L outperforms strong CTTA baselines with fewer parameters. In the case of w/o WAD, although TENT and SAR utilize fewer parameters, they suffer from severe forgetting at 10 rounds. Otherwise, our BECoTTA achieves +48.2%p, +7.7%p improvement than TENT and SAR respectively at the last round. In addition, our BECoTTA-S demonstrates 1.81%p gain using only $\sim$98fewer parameters (0.09M) than EcoTTA, while preserving the previous domain knowledge. Over CoTTA, all of our BECoTTA-S, M, and L achieve +1%p, +1.4%p and +2.1%p increased performance using 608$\times$, 86$\times$ and 17$\times$ reduced parameter, respectively. In the case of w/ WAD, our method surpasses other baselines that utilize fully updated source models even only updating MoDE layers. In particular, our BECoTTA-L shows a 16.8%p improvement over w/ WAD EcoTTA, which similarly updates only MetaNet as ours.
CDS-Easy (balanced weather shifts). As demonstrated in Tab. 2, BECoTTA achieves superior performance over other strong baselines. Compared within w/o WAD only, our BECoTTA-S outperforms EcoTTA (+5%p) and CoTTA (+0.1%p) by using only 2% and 0.1% number of parameters they used. Additionally, while BECoTTA-S uses a similar level of parameters (0.09M) as TENT and SAR, we demonstrate a +5%p performance increase compared to them. Ultimately, our BECoTTA succeeds in achieving +11.1%p higher performance than the source-only.
Continual Gradual Shifts (CGS).
@@ 4.3 Main Results
BECoTTA (Ours) - S 61.85 46.95 57.64 56.96 55.85 0.09M
                 + WAD 62.09 51.08 59.90 57.72 57.69 0.12M
BECoTTA (Ours) - M 60.49 46.20 58.24 57.45 55.60 0.63M
                 + WAD 64.04 53.25 60.66 58.55 59.13 0.77M
BECoTTA (Ours) - L 62.55 47.72 59.30 59.02 57.13 3.16M
                 + WAD 64.62 53.54 62.59 60.17 60.23 3.31M

### v3

@@ 3.4 Continual Test-time Adaptation Process
Model Initialization. Following recent trends in CTTA (Song et al., 2023; Niu et al., 2022; Lim et al., 2023; Choi et al., 2022a; Liu et al., 2021a; Adachi et al., 2022; Jung et al., 2023) and for a fair comparison, we perform a short pre-training for trainable parameters in models before deploying them to the CTTA problems. We initialize our BECoTTA with three different manners: (i) random, (ii) source-domain-only (w/o SDA, BECoTTA), and (iii) domain-augmented (w/ SDA, BECoTTA+) initialization. For (i) random initialization, we randomly initialize MoDE layer weights. For (ii) source-domain-only initialization, following EcoTTA (Song et al., 2023), we initialize MoDE layer weights using the source domain. Note that this initialization strategy regards the fair comparison in the CTTA literature. For (iii) domain-augmented initialization, we first build $D$ domains using SDA. Next, we introduce a Domain Discriminator (DD) as the auxiliary head. It consists of lightweight CNN layers and is trained to classify the pre-defined $D$ domains. This helps the model distinguish between different domains and classify each test-time image input accordingly during test-time adaptation on a sequence of unseen domains. Then, we update MoDE layers and DD for only a small number of epochs. The total initialization loss $L_{init}$ is formulated below. Except for the original cross-entropy loss $L_{seg}$ for semantic segmentation, we include the cross-entropy loss $L_{disc}$ for DD, and the domain-expert synergy loss $\Theta(D;A)$:
@@
Source-free CTTA with MoDE. Building upon the initialized MoDE, we deploy our BECoTTA to the continual target domains $\bm{X_{t}}$. Note that we do not access any source dataset after deployment, maintaining the source-free manner in the test time as other prior works. In the source-free TTA, we only activated MoDE layers to transfer the target domain knowledge efficiently. Utilizing the frozen DD trained at initialization, we obtain the pseudo-domain label $d$ for each domain-agnostic target image $x_{t}^{c}$. Afterward, according to $d$, we initially assign the domain-wise router and proceed with the aggregation of domain-adaptive experts. This approach ensures that our BECoTTA+ maintains input dependency, even within unseen test-time domains.
Table 1: Results on CDS-Hard (imbalanced weather & area shifts). We devise a novel scenario encompassing imbalanced weather and area shifts. We present performance results for both w/o SDA and w/ SDA across the overall baselines. We report $S$, $M$, and $L$ versions for our BECoTTA based on the number of parameters.
Round 1 10 $\Delta$ Parameter
Init Method Activated B-Clear A-Fog A-Night A-Snow B-Overcast Mean B-clear A-Fog A-Night A-Snow B-Overcast Mean
w/o SDA SAR (Niu et al., 2023) - 41.0 64.5 33.4 54.5 46.6 48.0 41.3 64.3 31.6 54.2 46.6 47.6 -0.4 0.02M
BECoTTA (S) MoDE 42.9 68.5 35.0 57.2 47.8 50.5 43.0 68.5 35.1 57.3 48.8 50.7 +0.1 0.09M
BECoTTA (M) MoDE 43.8 68.8 34.9 57.9 49.2 50.9 43.7 68.8 34.5 57.9 49.2 50.9 +0.0 0.63M
BECoTTA (L) MoDE 43.9 69.1 35.0 58.3 50.2 51.3 44.0 69.1 35.1 58.3 50.2 51.3 +0.0 3.16M
w/ SDA SAR (Niu et al., 2023) Full 43.6 68.6 44.5 59.1 48.7 52.9 43.4 67.4 42.2 58.1 47.6 51.9 -1.0 0.02M
BECoTTA+ (S) MoDE 44.1 69.5 40.1 56.8 49.1 51.9 44.0 69.4 40.1 56.9 49.1 51.9 +0.0 0.12M
BECoTTA+ (M) MoDE 45.6 70.8 42.6 59.6 50.8 53.9 45.6 70.7 42.5 59.5 50.8 53.9 +0.0 0.77M
BECoTTA+ (L) MoDE 45.7 71.4 43.7 59.6 50.5 54.2 45.7 71.3 43.7 59.6 50.6 54.2 +0.0 3.32M
Following preliminary works (Wang et al., 2020; Song et al., 2023; Niu et al., 2022), we adopt the entropy minimization using $H(\hat{y_{c}})=-\sum p(\hat{y_{c}})\cdot\log p(\hat{y_{c}})$. To avoid forgetting and error accumulation, we perform entropy filtering based on the confidence of the pseudo-labels. Therefore, the entropy-based loss $L_{tta}$ is as follows:
@@ 4 Experiments
We first introduce the datasets in Sec. 4.1, used for three continual segmentations, two classifications, and a domain generalization benchmark. Next, we describe the experimental setup in Sec. 4.2. Then, we provide our main results and analysis in Secs. 4.3 and 4.4, respectively. More detailed results are in the Appendix.
4.1 Datasets
@@
Continual Disjoint Shifts (CDS) benchmark. To reflect various domain shifts, we adopt balanced weather shifts (CDS-Easy) and imbalanced weather & area shifts (CDS-Hard) scenarios. For the CDS-Easy, we utilize the Cityscapes-ACDC setting used in previous work (Wang et al., 2022a): Cityscapes (Cordts et al., 2016) is used as the source domain, and ACDC (Sakaridis et al., 2021) as the target domain, consisting of four different weather types (fog, night, rain, snow). For the CDS-Hard, we propose a new imbalanced scenario considering both weather and geographical domain shifts. We also add clear and overcast weather from BDD-100k (Yu et al., 2020) to the existing target domain to mimic the real-world variety.
@@
Continual Gradual Shifts (CGS) benchmark. To construct gradually changing weather scenarios with blurry boundaries, we define sampling distributions per weather and perform uniform sampling. Next, we introduce four tasks containing blurred boundaries of weather, as illustrated in Fig. 3. The detailed process is in Appendix (Sec. D).
Classification benchmark. We additionally evaluate classification scenarios on CIFAR10-CIFAR10C (Hendrycks & Dietterich, 2019) and CIFAR100-CIFAR100C (Krizhevsky et al., 2009) with a non-ViT backbone.
@@
Domain Generalization (DG) benchmark. To demonstrate the versatility of BECoTTA, we conduct additional zero-shot experiments using the DG benchmark (Choi et al., 2021). This benchmark includes two large-scale real-world datasets (BDD-100k (Yu et al., 2020), Mapillary (Neuhold et al., 2017)) and two simulated datasets (GTAV (Richter et al., 2016), Synthia (Ros et al., 2016)).
@@
BN Stats Adapt (Nado et al., 2020a) - 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 52.0 0.09M
@@
BECoTTA (S) 71.3 41.1 62.4 59.8 71.3 41.1 62.4 59.8 71.4 41.1 62.4 59.8 58.6 0.09M
                                   + SDA 72.0 45.4 63.7 60.0 71.7 45.2 63.6 60.1 71.7 45.4 63.6 60.1 60.2 0.12M
BECoTTA (M) 72.3 42.0 63.5 60.1 72.4 41.9 63.5 60.2 72.3 41.9 63.6 60.2 59.5 2.15M
                                   + SDA 71.8 48.0 66.3 62.0 71.7 47.7 66.3 61.7 71.8 47.7 66.3 61.9 61.9 2.70M
BECoTTA (L) 71.5 42.6 63.2 59.1 71.5 42.6 63.2 59.1 71.5 42.5 63.2 59.1 59.1 11.31M
                                   + SDA 72.7 49.5 66.3 63.1 72.6 49.4 66.3 62.8 72.5 49.7 66.2 63.1 63.0 11.86M
4.2 Experimental Setting
Baselines. We compare our model with strong continual test-time adaptation methods including TENT (Wang et al., 2020), CoTTA (Wang et al., 2022a), SAR (Niu et al., 2023), EcoTTA (Song et al., 2023), VDP (Gan et al., 2023), DePT (Gao et al., 2022), TTN (Lim et al., 2023). More details are found in Appendix (Sec. A).
Evaluation metric. All of the semantic segmentation results are reported mIoU in %. For the overall scenarios, we repeat each task in 10 rounds (a few rounds are reported for visibility). Please refer to the Appendix (Sec. C) for the whole results. For the classification tasks, we report the classification error rate (%) following other baselines.
@@ 4.2 Experimental Setting
BECoTTA (S) 61.85 46.95 57.64 56.96 55.85 0.09M
                   + SDA 62.09 51.08 59.90 57.72 57.69 0.12M
BECoTTA (M) 60.49 46.20 58.24 57.45 55.60 0.63M
                   + SDA 64.04 53.25 60.66 58.55 59.13 0.77M
BECoTTA (L) 62.55 47.72 59.30 59.02 57.13 3.16M
                   + SDA 64.62 53.54 62.59 60.17 60.23 3.31M
Implementation details. BECoTTA has a flexible architecture design, so it provides multiple variants according to the selection of the expert rank ($dim$), the location of MoDE, the number of experts ($N$), and domain routers ($D$). Regarding $D$, we adopt D=1 for BECoTTA and D=4 for BECoTTA+. Regarding $dim$ and $N$, we categorize the results into three groups: S, M, and L. More specifically, we set four experts for S, and only inject MoDE into the last block of the encoder. Both M and L utilize six experts and inject MoDE into every block of the encoder. The difference between M and L is the $dim$ setting. More variants are found in Tab. 21.
For the CDS-Easy scenario, we leverage the pre-trained Segformer-B5 as our source model, aligning with CoTTA (Wang et al., 2022a). For other scenarios, we opt for Segformer-B2. Note that there is a difference between the two setups due to the size of Segformer, but we unify the Ours-S, M, and L architecture settings for all scenarios. To implement the initialization process, we warm up our architecture for 10 epochs like previous works (Song et al., 2023; Lim et al., 2023). For the classification task, we adopt the non-ViT backbones, WideResNet-28 for CIFAR10-CIFAR10C and WideResNet-40 for CIFAR100-CIFAR100C, for a fair comparison with other baselines. We provide further implementation details in Appendix (Sec. B).
Fairness with other baselines. We report both w/o SDA (i.e., BECoTTA) and w/ SDA (i.e., BECoTTA+) results for all experiments. In w/ SDA setup, we perform a slight initialization while activating full model parameters for the baselines (Wang et al., 2022a, 2020; Niu et al., 2023) that update full parameters or normalization layers only during CTTA. On the other hand, CTTA methods with parameter-efficient modules, such as EcoTTA and ours, perform initialization using SDA by updating these trainable modules only while freezing the pre-trained backbone weights.
Table 4: Classification error rate (%) for CIFAR100-to-CIFAR100C with severity level 5. Results are evaluated on WideResNet-40.
Method Gaus. Shot Impu. Defo. Glas. Moti. Zoom Snow Fros. Fog Brig. Cont. Elas. Pixe. Jpeg Avg. err
Source 80.1 77.0 76.4 59.9 77.6 64.2 59.3 64.8 71.3 78.3 48.1 83.4 65.8 80.4 59.2 69.7
tBN (Nado et al., 2020b) 45.9 45.6 48.2 33.6 47.9 34.5 34.1 40.3 40.4 47.1 31.7 39.7 42.7 39.2 45.6 41.1
Single do. TENT (Wang et al., 2020) 41.2 40.6 42.2 30.9 43.4 31.8 30.6 35.3 36.2 40.1 28.5 35.5 39.1 33.9 41.7 36.7
Continual TENT (Wang et al., 2020) 41.2 38.2 41.0 32.9 43.9 34.9 33.2 37.7 37.2 41.5 33.2 37.2 41.1 35.9 45.1 38.3
TTT++ (Liu et al., 2021b) 46.0 45.4 48.2 33.5 47.7 34.4 33.8 39.9 40.2 47.1 31.8 39.7 42.5 38.9 45.5 41.0
SWRNSP (Choi et al., 2022b) 42.4 40.9 42.7 30.6 43.9 31.7 31.3 36.1 36.2 41.5 28.7 34.1 39.2 33.6 41.3 36.6
NOTE (Gong et al., 2022) 50.9 47.4 49.0 37.3 49.6 37.3 37.0 41.3 39.9 47.0 35.2 34.7 45.2 40.9 49.9 42.8
EATA (Niu et al., 2022) 41.6 39.9 41.2 31.7 44.0 32.4 31.9 36.2 36.8 39.7 29.1 34.4 39.9 34.2 42.2 37.1
CoTTA (Wang et al., 2022a) 43.5 41.7 43.7 32.2 43.7 32.8 32.2 38.5 37.6 45.9 29.0 38.1 39.2 33.8 39.4 38.1
EcoTTA (Song et al., 2023) 42.7 39.6 42.4 31.4 42.9 31.9 30.8 35.1 34.8 40.7 28.1 35.0 37.5 32.1 40.5 36.4
BECoTTA (w/o SDA) 42.1 38.0 42.2 30.2 42.9 31.7 29.8 35.1 33.9 38.5 27.9 32.0 36.7 31.6 39.9 35.5
4.3 Main Results
CDS-Hard (imbalanced weather & area shifts). As shown in Tab. 1 and Fig. 1, all of our BECoTTA-S/M/L outperforms strong CTTA baselines with fewer parameters. In the case of w/o SDA, although TENT and SAR utilize fewer parameters, they suffer from severe forgetting at 10 rounds. Otherwise, our BECoTTA achieves +48.2%p, +7.7%p improvement than TENT and SAR respectively at the last round. In addition, our BECoTTA(S) demonstrates 1.81%p gain using only $\sim$98fewer parameters (0.09M) than EcoTTA, while preserving the previous domain knowledge. Over CoTTA, all of our BECoTTA(S), (M), and (L) achieve +1%p, +1.4%p and +2.1%p increased performance using 608$\times$, 86$\times$ and 17$\times$ reduced parameter, respectively. In the case of w/ SDA, our method surpasses other baselines that utilize fully updated source models even only updating MoDE layers. In particular, our BECoTTA+(L) shows a 16.8%p improvement over w/ SDA EcoTTA, which similarly updates only MetaNet as ours.
CDS-Easy (balanced weather shifts). As demonstrated in Tab. 2, BECoTTA achieves superior performance over other strong baselines. Compared within w/o SDA only, our BECoTTA(S) outperforms EcoTTA (+5%p) and CoTTA (+0.1%p) by using only 2% and 0.1% number of parameters they used. Additionally, while BECoTTA(S) uses a similar level of parameters (0.09M) as TENT and SAR, we demonstrate a +5%p performance increase compared to them. Ultimately, our BECoTTA succeeds in achieving +11.1%p higher performance than the source-only.